# Deep End-to-end Adaptive k-Space Sampling, Reconstruction, and Registration for Dynamic MRI

### v1

@@ 1 Introduction
Several traditional and DL-based methods exist for image registration in medical imaging, with the goal of aligning a moving image to a reference frame and estimating motion [24, 5, 10, 9, 40]. Within the scope of accelerated MRI, some approaches have combined registration with a reconstruction step by co-optimizing both processes [13, 41, 42], while others employ learned motion models [53].
@@
  • We present a detailed methodology covering all components of our approach, ensuring reproducibility for each module. Code will be made available.
  • We rigorously evaluate our framework under various configurations.
  • We validate our approach on a cardiac cine dataset and demonstrate its generalizability on out-of-distribution data (aorta dataset).
@@ 2.1 Dynamic MRI Reconstruction
$\mathbf{x}_{\cdot,\tau}:=\text{RSS}\circ\mathcal{F}^{-1}(\mathbf{y}_{\cdot,\cdot,\tau}),\quad\tau=1,\cdots,n_{t},$ (1)
where,
$\text{RSS}(\mathbf{w})=\big{(}\sum_{k=1}^{n_{c}}|\mathbf{w}_{\cdot,k}|^{2}\big{)}^{1/2}\in\mathbb{R}^{n},\quad\mathbf{w}\in\mathbb{C}^{n\times n_{c}}.$ (2)
@@
$\mathcal{C}_{\mathbf{S}_{\tau}}(\mathbf{z})\,=\,\left(\mathbf{S}_{\tau}^{1}\mathbf{z},\cdots,\mathbf{S}_{\tau}^{n_{c}}\mathbf{z}\right),\quad\text{for }\mathbf{z}\in\mathbb{C}^{n}.$ (5)
@@ 2.2 Deep Learning-based MRI Reconstruction
Deep learning methods have been widely applied to address accelerated MRI, bypassing the need for handcrafted priors and computationally intensive optimization techniques in CS. DL methods primarily fall into two categories: direct mapping and unrolled optimization approaches [11]. Direct mapping DL methods take undersampled images as input and learn to output reconstructed images, a straightforward approach that typically requires a substantial amount of training data. In contrast, unrolled optimization methods are inspired by compressed sensing techniques and leverage algorithms like gradient descent [17, 39, 60], primal-dual descent [1], conjugate gradient [2], and ADMM [58] to iteratively solve (6).
@@ 2.4 Deformable Image Registration
In this work we only employ a similarity measure as in (7) due to the absence of such segmentations within our datasets.
@@ 3.1.3 Reconstruction Model
$\mathcal{R}_{\mathbf{S}_{\tau}}(\mathbf{v}):=\sum_{k=1}^{n_{c}}(\mathbf{S}_{\tau}^{k})^{*}\mathbf{v}_{\cdot,k}\in\mathbb{C}^{n},\quad\text{for }\mathbf{v}\in\mathbb{C}^{n\times n_{c}}.$ (13)
Additional details on the reconstruction network are provided in Appendix 0.A.
@@ 3.3 Training Loss Function
With the end-to-end framework established, the primary objective is to optimize the registration task, specifically enhancing the accuracy of the deformation fields. Although the reconstruction task is secondary, we expect that improving the image reconstruction indirectly supports better motion estimation by providing high-fidelity inputs for alignment.
@@
$\begin{gathered}\mathcal{L}_{\text{rec}}=\mathcal{L}_{\text{ssim}}\left(|\hat{\mathbf{x}}_{\text{mov}}|,{\mathbf{x}}_{\text{mov}}\right)+\big{|}\big{|}|\hat{\mathbf{x}}_{\text{mov}}|-{\mathbf{x}}_{\text{mov}}\big{|}\big{|}_{1},\end{gathered}$ (16)
with $\hat{\mathbf{x}}_{\text{mov}}=\mathcal{V}_{\bm{\theta}}(\tilde{\mathbf{y}}_{\text{mov}}^{\mathbf{M}};\mathbf{S})$ representing the predicted moving image and $\mathbf{x}_{\text{mov}}=\text{RSS}\circ\mathcal{F}(\mathbf{y}_{\text{mov}})$ denoting the ground truth. Here, $\mathcal{L}_{\text{ssim}}:=\mathcal{L}_{\text{ssim2D}}+\mathcal{L}_{\text{ssim3D}}$ combines 2D and 3D structural similarity index measure (SSIM) losses [58], calculated over individual slices and across the sequence, respectively. For the registration loss $\mathcal{L}_{\text{reg}}$, we focus on aligning the registered reconstructed moving and reference images while ensuring a smooth deformation field. This includes a similarity term for image alignment and a regularization term to constrain the registration field:
@@
where $\mathbf{x}_{\text{reg}}=\big{\{}\mathcal{W}\big{(}\left|\hat{\mathbf{x}}_{\text{mov}}\right|_{\cdot,\tau},\bm{\phi}_{\tau}\big{)}\big{\}}_{\tau=1}^{n_{t}}\in\mathbb{R}^{n\times n_{t}}$ and $\mathbf{x}_{\text{ref}}=\text{RSS}\circ\mathcal{F}(\mathbf{y}_{\text{ref}})\in\mathbb{R}^{n}$. Note that $\left\{{\mathbf{x}}_{\text{ref}}\right\}_{t=1}^{n_{t}}\in\mathbb{R}^{n\times n_{t}}$ denotes ${\mathbf{x}}_{\text{ref}}$ repeated $n_{t}$ times. For $\mathcal{L}_{\text{sim}}$, we apply the same loss functions as in (16), and $\bm{\phi}=\mathcal{R}_{\bm{\psi}}(\big{[}|\hat{\mathbf{x}}_{\text{mov}}|,\mathbf{x}_{\text{ref}}\big{]})$ denotes the registration field prediction.
The smoothness term, inspired by a diffusion regularizer [12, 5], $\mathcal{L}_{\text{smooth}}$, aims to regularize the registration field by penalizing abrupt changes, approximating partial derivatives along each dimension through mean absolute finite differences:
$\mathcal{L}_{\text{smooth}}(\bm{\phi})=\frac{1}{2nn_{t}}\sum_{\tau}^{n_{t}}\sum_{\mathbf{p}}\left|\frac{\partial{\bm{\phi}_{\tau}}(\mathbf{p})}{\partial u_{1}}\right|+\left|\frac{\partial{\bm{\phi}_{\tau}}(\mathbf{p})}{\partial u_{2}}\right|,$ (18)
where $\frac{\partial{\bm{\phi}_{\tau}}(\mathbf{p})}{\partial u_{i}}$ denotes the approximated partial derivatives.
@@ 4.1 Datasets
We used the CMRxRecon 2023 cardiac cine dataset [49, 48], which comprises 472 scans with fully sampled, ECG-triggered multi-coil ($n_{c}=10$) $k$-space data, totaling 3,185 2D sequences, with the cardiac cycle segmented into 12 temporal frames. For external validation, we used the aorta dataset from CMRxRecon 2024 [50], also ECG-triggered with $n_{c}=10$ coils and 12-frame cycles. The inference subset comprises 111 scans and 1,332 sequences.
@@ 4.4 Evaluation
We assess estimated motion quality by evaluating the similarity of the registered moving images (warped reconstructed images using predicted deformation fields) to the reference image. We employ three image quality metrics: SSIM, PSNR, and NMSE. For brevity, definitions are provided in in Appendix 0.B. For each metric $m$, we averaged results across all phases of the sequence:
@@ 4.6 Ablation and Component Analysis
To assess our lightweight registration approach (concatenating moving and reference images along the channel dimension in a 2D U-Net), we compare it against DL-based alternatives:
@@
    A standard registration method - VoxelMorph [5]
@@
    Our proposed model using a larger U-Net
    and traditional motion estimation methods:
@@
See Appendix 0.B for the choice of hyperparameters.
@@
To investigate how task weighting impacts registration quality, we set the weighting parameters in (15) to $\alpha=1$, $\beta=3$. This allows us to examine the extent to which additional emphasis on the registration task influences the final motion estimation.
@@ 4.7 Experimental Results
We present average quantitative results for each experiment across both datasets (cine - seen during training, aorta - unseen), displayed as line graphs (function of acceleration factor) with ±$\pm 0.1$ standard deviations, along with average inference times.
Detailed results are also in tabular form in Appendix 0.C. Examples of qualitative results are provided in Figure 10, with further examples in Appendix 0.C.
@@
(a) Cine test dataset.
(b) Aorta dataset (not seen during training).
Figure 5: Comparison of registration performance across different methods.
Figure 5 presents the registration performance of our pipeline across various registration modules, evaluated on both datasets. Across all acceleration factors (4$\times$, 6$\times$, 8$\times$), our proposed lightweight model consistently outperforms traditional methods (OptFlow ILK, TV-$L_{1}$, DEMONS) and the two considered DL-based approaches (Voxelmorph, Transmorph) in all metrics on both datasets, indicating more accurate motion estimation. Notably, the larger variant of our lightweight model did not improve registration. Additionally, our pipeline with learned registration methods (including ours) achieves significantly lower inference times than non-DL-based methods.
@@
(a) Cine test dataset.
(b) Aorta dataset (not seen during training).
Figure 6: Impact of reconstruction model on registration results.
@@
(a) Cine test dataset.
(b) Aorta dataset (not seen during training).
Figure 7: Impact of phase-specific vs unified adaptive sampling with ACS and equispaced initialization on registration results.
@@
(a) Cine test dataset.
(b) Aorta dataset (not seen during training).
Figure 8: Impact of learned vs fixed non-adaptive sampling schemes on registration results.
@@
Although the reconstruction quality of the moving image is not a primary focus, we examine the effect of varying loss weights on both reconstruction and registration tasks in Figure 9. Results show that a higher registration weight ($\beta=3$) noticeably reduces reconstruction quality, yet registration accuracy improves across all metrics and undersampling rates.
@@
The results in Figure 9 further reveal that the joint training approach performs slightly better in terms of both, reconstruction and registration, than the decoupled method for equal weighting.
(a) Reconstruction performance.
(b) Registration performance.
Figure 9: Impact of joint vs decoupled training, and increased registration loss weight (cine dataset).
5 Discussion and Conclusion
@@
Finally, we compared joint versus decoupled training schemes. In the decoupled setup, task-specific losses were optimized by freezing the registration model weights during reconstruction loss computation, and vice versa. While joint optimization yielded only modest improvements over decoupled training, its value becomes clearer in our experiment with increased weighting on the registration loss. In this setting, prioritizing registration led to improved registration metrics despite a decline in reconstruction quality, highlighting how joint training allows flexible task prioritization. Future work could explore setting $\alpha=0$, focusing exclusively on registration to evaluate its standalone impact.
@@
Appendix 0.A Methods - Additional Information
0.A.1 Sensitivity Profile Network
@@
These initial sensitivity estimates, consistent with the methodology in [39], are further refined using a deep learning-based model. Specifically, a two-dimensional U-Net [35], denoted as $\mathcal{S}_{\bm{\sigma}}$, is employed to enhance the sensitivity profiles:
@@
0.A.2 Reconstruction Network - vSHARP
In our original pipeline, we employed the variable Splitting Half-quadratic ADMM algorithm for Reconstruction of inverse Problems (vSHARP) as the reconstruction network. vSHARP is an unrolled, physics-guided deep learning framework [58], which has demonstrated effectiveness in accelerated dynamic cardiac MRI reconstruction. Notably, it was among the winning solutions of the CMRxRecon 2023 and 2024 challenges [55, 57]. The vSHARP algorithm leverages the half-quadratic variable splitting technique for the optimization problem defined in (6), introducing an auxiliary variable $\mathbf{z}$ as follows:
@@ 0.A.2 Reconstruction Network - vSHARP
$\mathbf{x}^{t+1}=\operatorname*{argmin}_{\mathbf{x}^{{}^{\prime}}\,\in\,\mathbb{C}^{n\times n_{t}}}\frac{1}{2}\sum_{\tau=1}^{n_{t}}\left|\left|\mathcal{T}_{\mathbf{M}_{\tau},\mathbf{C}_{\tau}}(\mathbf{x}_{\cdot,\tau}^{{}^{\prime}})-\tilde{\mathbf{y}}_{\cdot,\cdot,\tau}^{\mathbf{M}}\right|\right|_{2}^{2}+\lambda^{t+1}\left|\left|\mathbf{x}^{{}^{\prime}}-\mathbf{z}^{t+1}+\frac{\mathbf{m}^{t}}{{\lambda}^{t+1}}\right|\right|_{2}^{2},$ (A5b)
@@
In (A5a), $\mathcal{D}_{\bm{\theta}^{t+1}}$ denotes a convolutional DL-based image denoiser with trainable parameters $\bm{\theta}^{t+1}$, and $\lambda^{t+1}$ is a trainable learning rate. At each iteration, $\mathcal{D}{\bm{\theta}^{t+1}}$ processes the previous estimates of the three variables, yielding an updated estimate for $\mathbf{z}$. Equation A5b is numerically optimized using a differentiable gradient descent scheme unrolled over $T_{\mathbf{x}}$ iterations. The update in (A5c) involves a straightforward calculation. The initial estimates for $\mathbf{x}$ and $\mathbf{z}$ are defined as follows:
@@
Appendix 0.B Experiments - Additional Information
0.B.1 Datasets

### v2

@@ 1 Introduction
Several traditional and DL-based methods exist for image registration in medical imaging, with the goal of aligning a moving image to a reference frame and estimating motion [24, 5, 10, 9, 40]. Within the scope of accelerated MRI, some approaches have combined registration with a reconstruction step by co-optimizing both processes [13, 41, 42], while others employ learned motion models that bypass direct reconstruction [54].
@@
  • We present a detailed methodology covering all components of our approach, ensuring reproducibility for each module. We will also make our code available.
  • We rigorously evaluate our framework under various configurations. We validate our approach on a cardiac cine dataset and demonstrate its generalizability on out-of-distribution data (aorta dataset).
@@ 2.1 Dynamic MRI Reconstruction
$\begin{gathered}\mathbf{x}_{\cdot,\tau}:=\text{RSS}\circ\mathcal{F}^{-1}(\mathbf{y}_{\cdot,\cdot,\tau}),\quad\tau=1,\cdots,n_{t},\\ \text{RSS}(\mathbf{w})=\big{(}\sum_{k=1}^{n_{c}}|\mathbf{w}_{\cdot,k}|^{2}\big{)}^{1/2}\in\mathbb{R}^{n},\quad\mathbf{w}\in\mathbb{C}^{n\times n_{c}}.\end{gathered}$ (1)
@@
$\mathcal{C}_{\mathbf{S}_{\tau}}(\mathbf{z})\,=\,\left(\mathbf{S}_{\tau}^{1}\mathbf{z},\cdots,\mathbf{S}_{\tau}^{n_{c}}\mathbf{z}\right),\quad\text{for}\quad\mathbf{z}\in\mathbb{C}^{n}.$ (4)
@@ 2.2 Deep Learning-based MRI Reconstruction
Deep learning methods have been widely applied to address accelerated MRI, bypassing the need for handcrafted priors and computationally intensive optimization techniques in CS. DL methods primarily fall into two categories: direct mapping and unrolled optimization approaches [11]. Direct mapping DL methods take undersampled images as input and learn to output reconstructed images, a straightforward approach that typically requires a substantial amount of training data. In contrast, unrolled optimization methods are inspired by CS and leverage algorithms like gradient descent [17, 39, 59], primal-dual descent [1], conjugate gradient [2], and ADMM [57] to iteratively solve (5).
@@ 3.1.1 Sensitivity Profiles Estimation
The refined sensitivity profiles are subsequently normalized to satisfy
@@ 3.1.3 Reconstruction Model
$\mathcal{R}_{\mathbf{S}_{\tau}}(\mathbf{v}):=\sum_{k=1}^{n_{c}}(\mathbf{S}_{\tau}^{k})^{*}\mathbf{v}_{\cdot,k}\in\mathbb{C}^{n},\quad\text{for}\quad\mathbf{v}\in\mathbb{C}^{n\times n_{c}}.$ (12)
Additional details on the reconstruction network are provided in Appendix A.
@@ 3.3 Training Loss Function
With the end-to-end framework established, the primary objective is to optimize the registration task, enhancing the accuracy of the deformation fields. Although the reconstruction task is secondary, we expect that improving the image reconstruction indirectly supports better motion estimation by providing high-fidelity inputs for alignment.
@@
$\begin{gathered}\mathcal{L}_{\text{rec}}=\mathcal{L}_{\text{ssim}}\left(|\hat{\mathbf{x}}_{\text{mov}}|,{\mathbf{x}}_{\text{mov}}\right)+\mathcal{L}_{1}(|\hat{\mathbf{x}}_{\text{mov}}|,{\mathbf{x}}_{\text{mov}}),\end{gathered}$ (15)
with $\hat{\mathbf{x}}_{\text{mov}}=\mathcal{V}_{\bm{\theta}}(\tilde{\mathbf{y}}_{\text{mov}}^{\mathbf{M}};\mathbf{S})$ representing the predicted moving image and $\mathbf{x}_{\text{mov}}=\text{RSS}\circ\mathcal{F}(\mathbf{y}_{\text{mov}})$ denoting the ground truth. Here, $\mathcal{L}_{\text{ssim}}:=\mathcal{L}_{\text{ssim2D}}+\mathcal{L}_{\text{ssim3D}}$ combines 2D and 3D structural similarity index measure (SSIM) losses [57], calculated over individual slices and across the sequence, respectively.
For the registration loss $\mathcal{L}_{\text{reg}}$, we focus on aligning the registered reconstructed moving and reference images while ensuring a smooth deformation field. This includes a similarity term for image alignment and a smoothness term to constrain the registration field:
@@
where $\mathbf{x}_{\text{reg}}=\big{\{}\mathcal{W}\big{(}\left|\hat{\mathbf{x}}_{\text{mov}}\right|_{\cdot,\tau},\bm{\phi}_{\tau}\big{)}\big{\}}_{\tau=1}^{n_{t}}\in\mathbb{R}^{n\times n_{t}}$ and $\mathbf{x}_{\text{ref}}=\text{RSS}\circ\mathcal{F}(\mathbf{y}_{\text{ref}})\in\mathbb{R}^{n}$. Note that $\left\{{\mathbf{x}}_{\text{ref}}\right\}_{t=1}^{n_{t}}\in\mathbb{R}^{n\times n_{t}}$ denotes ${\mathbf{x}}_{\text{ref}}$ repeated $n_{t}$ times. For $\mathcal{L}_{\text{sim}}$, we apply the same loss functions as in (15), and $\bm{\phi}=\mathcal{R}_{\bm{\psi}}(\big{[}|\hat{\mathbf{x}}_{\text{mov}}|,\mathbf{x}_{\text{ref}}\big{]})$ denotes the registration field prediction. The smoothness term, inspired by diffusion regularization [12, 5], penalizes abrupt changes in $\bm{\phi}$:
$\mathcal{L}_{\text{smooth}}(\bm{\phi})=\frac{1}{2nn_{t}}\sum_{\tau}^{n_{t}}\sum_{\mathbf{p}}\left|\frac{\partial{\bm{\phi}_{\tau}}(\mathbf{p})}{\partial u_{1}}\right|+\left|\frac{\partial{\bm{\phi}_{\tau}}(\mathbf{p})}{\partial u_{2}}\right|.$ (17)
@@ 4.1 Datasets
We used the CMRxRecon 2023 cardiac cine dataset [49, 48], which comprises 472 scans with fully sampled, ECG-triggered multi-coil ($n_{c}=10$) $k$-space data, totaling 3,185 2D sequences, with the cardiac cycle segmented into 12 temporal frames. For external validation, we used the aorta dataset from CMRxRecon 2024 [51, 50], also ECG-triggered with $n_{c}=10$ coils and 12-frame cycles. The inference subset comprises 111 scans and 1,332 sequences.
@@ 4.4 Evaluation
We assess estimated motion quality by evaluating the similarity of the registered moving images (warped reconstructed images using predicted deformation fields) to the reference image. We employ three image quality metrics: SSIM, PSNR, and NMSE. For brevity, definitions are provided in in Appendix B. For each metric $m$, we averaged results across all phases of the sequence:
@@ 4.6 Ablation and Component Analysis
To assess our lightweight registration approach (concatenating moving and reference images along the channel dimension in a 2D U-Net), we compare it against DL-based alternatives and traditional motion estimation methods:
@@
    A standard registration method - VoxelMorph [5], using a 2D U-Net with four scales (16, 32, 64, 128 filters).
@@
See Appendix B for the choice of hyperparameters.
@@
To investigate how task weighting impacts registration quality, we test varying weighting parameters in (14). This allows us to examine the extent to which additional emphasis on the registration task influences the final motion estimation.
@@
Choice of Loss Function
We test two variants:
  1. 1.
    using only $\mathcal{L}_{\text{reg}}=\mathcal{L}_{\text{sim}}=\mathcal{L}_{\text{ssim2D}}$,
  2. 2.
    using only $\mathcal{L}_{\text{reg}}=\mathcal{L}_{\text{sim}}=\mathcal{L}_{1}$.
@@ 4.7 Experimental Results
We present average quantitative results for each experiment across both datasets (cine - seen during training, aorta - unseen), displayed as line graphs (function of acceleration factor) with ±$\pm 0.1$ standard deviations, along with average inference times (s). Detailed results are also in tabular form in Appendix C. Examples of qualitative results are provided in Figure 11, with further examples in Appendix C.
Figure 5: Comparison of registration performance across different methods. Top: Cine. Bottom: Aorta (not seen during training).
Figure 6: Impact of reconstruction model on registration results. Top: Cine. Bottom: Aorta (not seen during training).
Figure 7: Impact of phase-specific vs unified adaptive sampling with ACS and equispaced initialization on registration results. Top: Cine. Bottom: Aorta (not seen during training).
Figure 8: Impact of learned vs fixed non-adaptive sampling schemes on registration results. Top: Cine. Bottom: Aorta (not seen during training).
Figure 9: Impact of joint (J) vs decoupled (D) training, and varying loss weights (cine dataset). Top: Reconstruction performance. Bottom: Registration performance.
Figure 10: Impact of loss choice (cine dataset). Top: Reconstruction performance. Bottom: Registration performance.
@@
Figure 5 presents the registration performance of our pipeline across various registration modules, evaluated on both datasets. Across all acceleration factors (4$\times$, 6$\times$, 8$\times$), our proposed lightweight model consistently outperforms traditional methods (OptFlow ILK, TV-$L_{1}$, DEMONS) and the two considered DL-based approaches (Voxelmorph, Transmorph) in all metrics on both datasets, indicating more accurate motion estimation. Additionally, our pipeline with learned registration methods (including ours) achieves significantly lower inference times than non-DL-based methods.
@@
Figure 9 illustrates the impact of varying loss weights on each task. Increasing the registration weight ($\alpha<\beta$) improves registration but degrades reconstruction. However, analysis of motion estimates and images (Appendix C, Figures C7 and C8) indicates that when $\alpha<\beta$, the network may approximate the reference rather than reconstructing the moving image accurately. Conversely, when $\alpha>\beta$, reconstruction quality improves at the cost of motion estimation. Among all weighting schemes, equal weighting preserves motion field accuracy while maintaining reconstruction quality.
@@
Figure 9 further contrasts joint vs. decoupled training. While joint optimization slightly reduces reconstruction performance, it significantly improves registration quality.
Choice of Loss Function
Results are shown in Figure 10. Both $\mathcal{L}_{\text{ssim}}$ and $\mathcal{L}_{1}$ are important for maintaining higher registration accuracy.
@@ 5 Discussion and Conclusion
In addition, we compared joint versus decoupled training schemes. In the decoupled setup, task-specific losses were optimized by freezing the registration model weights during reconstruction loss computation, and vice versa, with joint optimization yielding significant registration improvements over decoupled training, with minimal reconstruction sacrifice. This suggests that joint optimization fosters better motion estimation by aligning reconstruction with registration objectives, whereas decoupled training isolates the tasks at the expense of registration quality. However, when prioritizing registration weight, this led to improved registration as the reconstruction model was tricked into producing the reference image (as a result of the joint optimization). Conversely, prioritizing reconstruction weight, can hinder registration performance. These suggest that balanced weighting yields more reliable motion estimation.
Finally, we analyzed the impact of different components in $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{sim}}$, finding that incorporating all components led to the best overall performance.
@@
Acknowledgments
This work was supported by institutional grants of the Dutch Cancer Society and of the Dutch Ministry of Health, Welfare and Sport. The authors would like to acknowledge the Research High Performance Computing (RHPC) facility of the Netherlands Cancer Institute (NKI).
Supplementary Material
Appendix A Methods - Additional Information
A.1 Sensitivity Profile Network
@@
These initial sensitivity estimates, consistent with the methodology in [39], are further refined using a deep learning-based model. Specifically, a 2D U-Net [35], denoted as $\mathcal{S}_{\bm{\sigma}}$, is employed to enhance the sensitivity profiles:
@@
A.2 Reconstruction Network - vSHARP
In our original pipeline, we utilized the variable Splitting Half-quadratic ADMM algorithm for Reconstruction of inverse Problems (vSHARP) as the reconstruction network. vSHARP is an unrolled, physics-guided deep learning method [57] that has been successfully applied in accelerated dynamic cardiac MRI reconstruction, including among the winning solutions of the CMRxRecon 2023 and 2024 challenges [30, 50]. The vSHARP algorithm leverages the half-quadratic variable splitting technique for the optimization problem defined in (6), introducing an auxiliary variable $\mathbf{z}$ as follows:
@@ A.2 Reconstruction Network - vSHARP
$\mathbf{x}^{t+1}=\operatorname*{argmin}_{\mathbf{x}^{{}^{\prime}}\,\in\,\mathbb{C}^{n\times n_{t}}}\frac{1}{2}\sum_{\tau=1}^{n_{t}}\left|\left|\mathcal{T}_{\mathbf{M}_{\tau},\mathbf{C}_{\tau}}(\mathbf{x}_{\cdot,\tau}^{{}^{\prime}})-\tilde{\mathbf{y}}_{\cdot,\cdot,\tau}^{\mathbf{M}}\right|\right|_{2}^{2}$
$+\lambda^{t+1}\left|\left|\mathbf{x}^{{}^{\prime}}-\mathbf{z}^{t+1}+\frac{\mathbf{m}^{t}}{{\lambda}^{t+1}}\right|\right|_{2}^{2},$ (A5b)
@@
In (A5a), $\mathcal{D}_{\bm{\theta}^{t+1}}$ denotes a convolutional DL-based image denoiser with trainable parameters $\bm{\theta}^{t+1}$, and $\lambda^{t+1}$ is a trainable learning rate. At each iteration, $\mathcal{D}_{\bm{\theta}^{t+1}}$ processes the previous estimates of the three variables, yielding an updated estimate for $\mathbf{z}$. Equation A5b is numerically optimized using a differentiable gradient descent scheme unrolled over $T_{\mathbf{x}}$ iterations. The update in (A5c) involves a straightforward calculation. The initial estimates for $\mathbf{x}$ and $\mathbf{z}$ are defined as follows:
@@
Appendix B Experiments - Additional Information
B.1 Datasets